jigglypuff314 (jigglypuff314):

Calculus help please? Find the point on the parabola y^2 = 2x that is closest to the point (1, 4).

jigglypuff314 (jigglypuff314):

so I have to take the derivative of the distance formula right? which would be http://www.wolframalpha.com/input/?i=root%28+%28x-1%29%5E2+%2B+%28+%282x%29%5E%281%2F2%29+-+4%29%5E2+%29

jigglypuff314 (jigglypuff314):

but I was supposed to take the derivative by hand... but I had a brain fart... HELP?!

OpenStudy (anonymous):

y^2=2x => x=(1/2)y^2 any point on the parabola has the form (x, y), that is (0.5y^2, y) We may as well minimise the square of the distance, as it is easier to calculate. We can do this because distance is positive, suppose the distance be d, then: d^2 = (0.5y^2 - 1)^2 + (y - 4)^2 which is the formula for distance. Now expand the right-hand side. d^2 = 0.25y^4 -y^2 + 1 + y^2 - 8y +16 d^2 = 0.25y^4 - 8y + 17 minimise by setting the derivative equal to zero. y^3 - 8 = 0 So y =2 substitute this back into your formula for distance d = sqrt[ (0.5y^2 - 1)^2 + (y - 4)^2 ] d=sqrt[ 1 + 4 ] = sqrt(5). so sqrt(5) is the final result xD

OpenStudy (anonymous):

I hope I'm not wrong :P I suck at calculus xDD

jigglypuff314 (jigglypuff314):

thanks :) can't be worse than me ;) got a 62% on this test :P

OpenStudy (anonymous):

xDD I had 70% marks in my multi variable calculus & 75% in calculus 1

jigglypuff314 (jigglypuff314):

erm @annas how did ya get from d^2 = 0.25y^4 - 8y + 17 to y^3 - 8 = 0 wasn't there a d^2 on the left? :3 so would you derive implicitly or something?

OpenStudy (anonymous):

hmm I've no idea lol I just remember my instructor told me this ^_^ lol told you I'm pretty bad at calculus >.>

jigglypuff314 (jigglypuff314):

mk then :P

OpenStudy (anonymous):

lol wait I've called the expert ^_^

jigglypuff314 (jigglypuff314):

nevermind, it still works ;)

OpenStudy (anonymous):

okie dokie ^_^

jigglypuff314 (jigglypuff314):

@annas part b: verify that the tangent to the parabola y^2 = 2x at your answer is perpendicular to the line connecting your answer to the point (1, 4)

OpenStudy (anonymous):

@jigglypuff314 I dont remember how to verify but try the link I think they explained well. http://www.dummies.com/how-to/content/how-to-find-the-tangent-lines-of-a-parabola-that-p.html